# ZMM-TTS: Zero-shot Multilingual and Multispeaker Speech Synthesis Conditioned on Self-supervised Discrete Speech Representations

### v1

@@ Abstract
Neural text-to-speech (TTS) has achieved human-like synthetic speech for single-speaker, single-language synthesis. Multilingual TTS systems are limited to resource-rich languages due to the lack of large paired text and studio-quality audio data. In most cases, TTS systems are built using a single speaker’s voice. However, there is growing interest in developing systems that can synthesize voices for new speakers using only a few seconds of their speech. This paper presents ZMM-TTS, a multilingual and multispeaker framework utilizing quantized latent speech representations from a large-scale, pre-trained, self-supervised model. Our paper is the first to incorporate the representations from text-based and speech-based self-supervised learning models into multilingual speech synthesis tasks. We conducted comprehensive subjective and objective evaluations through a series of experiments. Our model has been proven effective in terms of speech naturalness and similarity for both seen and unseen speakers in six high-resource languages. We also tested the efficiency of our method on two hypothetical low-resource languages. The results are promising, indicating that our proposed approach can synthesize audio that is intelligible and has a high degree of similarity to the target speaker’s voice, even without any training data for the new, unseen language.
@@ I INTRODUCTION
Text-to-speech (TTS) technology, which converts a text string into a speech waveform[1, 2, 3, 4], has observed significant advancements driven by deep learning. Advances in this field have been beneficial for a variety of uses including audio-book narration, news readers, conversational assistants and engaging user experiences in virtual worlds. However, this success is dependent on a rich resource of high-quality data, and thus data requirements have become a bottleneck of both research and the implementation of neural TTS[5]. Preparing data resources and building TTS systems for a specific language can be costly, let alone repeating this for thousands of languages around the world. In this paper, our aim is to build a unified model for a multilingual and multispeaker system, opening possibilities for unseen speakers and unseen language adaptations in low-resource scenarios.
@@
The expansion of the coverage of the world’s languages by TTS systems has attracted significant attention from both academia and industry[6, 7, 8]. Phonetically-based speech processing systems often require pronunciation dictionaries for mapping words to a sequence of phonetic units. Recent studies have demonstrated that pre-trained models for phoneme representations, such as PnG BERT [9], Mixed-Phoneme BERT [10] and Phoneme-level BERT [11], can be beneficial for advanced TTS systems. Motivated by the impressive cross-lingual transferability of multilingual language models, some work [5, 12] perform masked-language model (MLM) pre-training with multilingual text-only data for TTS. Considering the abundant resources of textual data in the real world, which are usually more readily accessible compared to audio, it is worthwhile to investigate the utilization of language models pre-trained on large text resources for phoneme representations in multilingual synthesis tasks.
@@
To synthesize speech with a target speaker’s timbre, end-to-end models usually learn a speaker embedding table in training that enables speaker selection during inference. Furthermore, several of these models enabled zero-shot speaker synthesis using a speaker vector extracted from a short audio clip[13, 14]. Despite recent advances, the similarity gap between seen and unseen speakers remains an open research question, and training these models still requires a significant amount of speakers, posing challenges in developing high-quality models for low-resource languages[8]. Furthermore, in real-world data, most of the attributes that represent speech are interrelated and difficult to disentangle. To control a multilingual and multispeaker system, it is crucial to disentangle the speaker and language[6]. Efforts using adversarial learning or consistency loss [6, 8] have been made to mitigate the performance degradation resulting from this entanglement.
@@
However, these state-of-the-art multilingual and multispeaker systems mainly rely on a large amount of training data, which is not always available for every language. The intermediate features used in these studies are often Mel spectrograms, which have a high correlation in time and frequency, making it difficult to disentangle speaker-dependent information. Fortunately, self-supervised learning (SSL) speech representations [15, 16, 17] have been shown to be useful for speech processing tasks, such as speech recognition [18, 19], speech reconstruction [20], and voice conversion [21]. The learned representations are used as input for a supervised model, which is often fine-tuned to improve task performance or reduce the labeled data required. Recently, some TTS models[22, 23] have started using discrete vector quantized speech representations as intermediate features instead of traditional Mel spectrograms for prediction. As a result, the quantized output has less speaker-dependent information than the Mel spectrograms[23]. While a designed fine-tuning protocol for the learned representations indeed improves ASR performance [24], in low-resource settings, relatively little attention has been paid to the study of TTS models using learned SSL representations.
@@
In this paper, we propose a unified zero-shot multilingual multispeaker TTS framework called ZMM-TTS that leverages discrete speech representations from a multilingual speech-based SSL model. ZMM-TTS consists of a text-to-discrete speech representations module (txt2vec) and a representations-to-waveform module (vec2wav). Regarding the txt2vec module, we adopt a standard end-to-end TTS architecture comprising token embeddings, an encoder, and a decoder to predict discrete representations. To utilize the strong cross-lingual transferability of multilingual language models, we adopt a pre-trained large-scale multilingual language model for phoneme representations. For the vec2wav module, an additional multi-stage and multi-head VQ (Vector Quantization) model was adopted to preserve the discrete information at different time resolutions.
@@
  • We propose ZMM-TTS, a multilingual and multispeaker framework using discrete audio representations from a large-scale multilingual pre-trained self-supervised model as an intermediate representation to replace the traditional Mel spectrograms.
  • We investigate the impact of various input representations on multilingual synthesis tasks that use SSL discrete representations. Our research is the first to combine phoneme representations from a pre-trained text-based multilingual language model and speech-based SSL representations in a speech synthesis task.
@@
  • We also verified the effectiveness of the ZMM-TTS in low-resource and zero-shot scenarios on two hypothetical low-resource languages. This observation is quite promising, as it implies that our proposed approach can synthesize understandable audio with high speaker similarity, even when no specific training data are available for the previously unseen language. Note that “unseen” in this context refers to the absence of dedicated training data for ZMM-TTS, while the unseen language is still included in the training data of both the pre-trained language model and self-supervised model.
We encourage the reader to listen to our samples, which can be found at https://gongchenghhu.github.io/TASLP-demo/.
@@ II RELATED WORK
This section reviews related work about zero-shot and multilingual speech synthesis. We also review related studies on speech synthesis with the aid of SSL representations especially in the multilingual case.
@@ II-A1 Zero-shot multispeaker TTS
The concept of zero-shot multispeaker TTS was initially introduced in [25] and the main idea is to employ embeddings from an external speaker encoder as a reference signal. Subsequent research by [14] aimed to reduce the quality gap between seen and unseen speakers by utilizing more informative embeddings. [26] further explored the utilization of attentive speaker embeddings for the purpose of encoding general speaking styles, while [27] proposed a speaker-conditional architecture and investigated a flow-based decoder capable of generating the voice of unseen speakers. Despite these efforts, the task of zero-shot multispeaker TTS is still not fully solved. Moreover, capturing a wide range of voice properties requires a substantial amount of high-quality data from various speakers.
Regarding multispeaker multilingual TTS, a primary idea is to train the decoder by incorporating learnable language and speaker embeddings[6]. More specifically, the majority of speech attributes are interrelated and difficult to disentangle. Attempts have been made to mitigate the decline in multi-language multispeaker performance caused by the speaker and language entanglement, especially in cross-lingual scenarios. [6] incorporated adversarial domain training, enabling them to transfer different voices between languages. [28] proposed the use of mutual information minimization to keep speaker consistency in cross-lingual synthesis. [29] utilized joint training incorporating a speaker classifier to enhance speaker similarity.
@@ II-A2 System structure
Previous multilingual TTS models [6, 7] are primarily built using Tacotron [1, 3]. When synthesizing speech, Tacotron-based models often repeat or skip words due to an autoregressive approach that uses attention to align input text and target speech. Instead of autoregressive iteration, several non-autoregressive [30, 31, 2, 32] models adopt Transformer encoder and decoder blocks to generate Mel spectrograms in parallel. For example, [33] create a multilingual speech synthesis system including a Transformer-based acoustic predictor and a WaveNet [34] neural vocoder. [35] presented a multilingual, multiaccented, multispeaker speech synthesis model based on RADTTS [36] which is a parallel flow-based generative model. Recently, several fully end-to-end multilingual speech synthesis systems [37, 8] were built on VITS[38]. VITS could produce natural-sounding speech at a high synthesis speed. Moreover, it requires training a single model, so there is no need for an independent vocoder. Among these fully end-to-end systems, YourTTS [8], which aims at zero-shot multispeaker TTS and zero-shot voice conversion, presented promising results for a few language combinations. However, it had limited success in transferring to languages with few speakers and used a curriculum learning approach, making training cumbersome.
@@ II-B1 Characters or Phonemes
Traditionally, end-to-end text-to-speech (TTS) models have utilized input representations in the form of characters or phonemes[1, 2]. Employing each character as the default input for end-to-end TTS models requires the model to implicitly learn how to convert graphemes to phonemes as part of the synthesis task. Incorporating phoneme input simplifies the TTS task, eliminating the need for the model to learn complicated grapheme-to-phoneme rules for languages such as English. Furthermore, language-independent pronunciation information can be obtained through language expertise, such as making a unified set of phonemes, often derived from the International Phonetic Alphabet (IPA) [5, 39].
II-B2 Pre-trained phoneme representations
Over the past decade, self-supervised pre-training on extensive text corpora, employing language model (LM) or MLM objectives, has demonstrated remarkable success in the field of TTS[9, 10, 11]. Toward multilingual tasks, [12] proposed XPhoneBERT which is the first pre-trained multilingual model for phoneme representations for TTS. Based on the results of monolingual experiments, it has been found that using XPhoneBERT as a phoneme encoder for input significantly enhances the performance of a powerful neural TTS model in terms of naturalness and prosody. It also helps to produce high-quality speech under low-resource training data. [5] conducted pre-training on a multilingual text-only dataset using MLM. They then trained this model in a supervised manner on paired text and audio data, while keeping the language-aware embedding layer frozen. This approach allows the model to synthesize languages that are not present in the paired data but exist in the text-only data.
Alternative approaches adopt UTF-8 bytes [40] or phonological features (PFs) [41] as input representations. UTF-8 byte representations encode typographic rather than phonological information. The model doesn’t learn unseen byte combinations, requiring retraining when enrolling a new language. As for PFs, IPA symbols can be converted to PFs according to certain aspects of their articulation, as implemented in Epitran [42]. Although this conversion is based on an explicit database of IPA symbols and diacritics, it can be considered universally applicable to any language for which IPA transcription is available. However, there are no pre-trained PF encoders similar to XPhoneBERT currently available. The effect of different input representations is still not clear in multilingual speech synthesis with SSL representations. Taking into account the complexity of the model and the support for new languages and low-resource scenarios, we evaluated the effects of using characters, IPA, and XPhoneBERT as input for multilingual TTS.
@@ II-C SSL Representations for speech synthesis
Well-designed proxy tasks enable SSL to explore the characteristics of unlabeled data and enhance the model’s representation ability. The derived representations of the SSL model are found to have useful information for many downstream tasks[43, 44, 45, 46, 47]. Wav2vec 2.0 [15] is the most popular pre-trained model that uses SSL on a large amount of unlabeled speech data. Wav2vec 2.0 has demonstrated strong representation abilities in ASR tasks and achieved outstanding results. A multilingual wav2vec 2.0, denoted as XLSR-53, was also proposed in [48]. The results showed a significant improvement over monolingual wav2vec 2.0, particularly in languages with limited resources.
@@
For SSL in TTS downstream tasks, [44] introduced WavThruVec, a two-stage architecture that employed high-dimensional wav2vec 2.0 embeddings as an intermediate speech representation. Using wav2vec 2.0 acoustic representations of synthesized audio, [47] improved prosody modeling by incorporating prosodic characteristics from neighboring utterances. [23] demonstrated that the VQ features in wav2vec 2.0 contain significantly less speaker information than the Mel spectrograms, and proposed DSE-TTS, a cross-lingual TTS that encoded timbre-related and linguistic-related speaker information separately. In the paper [49], the author proposed extracting unsupervised phonetic representations (UPR) from wav2vec 2.0 for multilingual speech synthesis when target language pronunciation dictionaries are unavailable. In the work of [50], the author built a text-to-speech synthesis system for the Scottish Gaelic language based on self-supervised discrete acoustic unit sequences.
@@
However, integrating SSL-based representations into TTS in low-resource settings is still a novel concept, and it is not clear which type of input representation is most suitable for TTS. Furthermore, these methods mostly use monolingual speech representations rather than multilingual speech representations, which may benefit low-resource scenarios.

### v2

@@ Abstract
Neural text-to-speech (TTS) has achieved human-like synthetic speech for single-speaker, single-language synthesis. Multilingual TTS systems are limited to resource-rich languages due to the lack of large paired text and studio-quality audio data. TTS systems are typically built using a single speaker’s voice, but there is growing interest in developing systems that can synthesize voices for new speakers using only a few seconds of their speech. This paper presents ZMM-TTS, a multilingual and multispeaker framework utilizing quantized latent speech representations from a large-scale, pre-trained, self-supervised model. Our paper combines text-based and speech-based self-supervised learning models for multilingual speech synthesis. Our proposed model has zero-shot generalization ability not only for unseen speakers but also for unseen languages. We have conducted comprehensive subjective and objective evaluations through a series of experiments. Our model has proven effective in terms of speech naturalness and similarity for both seen and unseen speakers in six high-resource languages. We also tested the efficiency of our method on two hypothetically low-resource languages. The results are promising, indicating that our proposed approach can synthesize audio that is intelligible and has a high degree of similarity to the target speaker’s voice, even without any training data for the new, unseen language.
@@ I INTRODUCTION
Text-to-speech (TTS) technology, which converts a text string into a speech waveform[1, 2, 3, 4], has received significant advancements driven by deep learning. Advances in this field have been beneficial for a variety of uses including audio-book narration, news readers, conversational assistants and engaging user experiences in virtual worlds. However, this success is dependent on a rich resource of high-quality data, and thus data requirements have become a bottleneck of both research and the implementation of neural TTS[5]. Preparing data resources and building TTS systems for a specific language can be costly, let alone repeating this for thousands of languages around the world. In this paper, our aim is to build a unified model for a multilingual and multispeaker system, opening possibilities for unseen speakers and unseen language adaptations in low-resource scenarios.
@@
The expansion of the TTS systems’ coverage of the world’s languages has attracted significant attention from both academia and industry[6, 7, 8]. Phonetically-based speech processing systems often require pronunciation dictionaries to map words to a sequence of phonetic units. Recent studies have demonstrated that pre-trained models for phonetic representations, such as PnG BERT [9], Mixed-Phoneme BERT [10] and Phoneme-level BERT [11], can be beneficial for advanced TTS systems. Motivated by the impressive cross-lingual transferability of multilingual language models, several works [5, 12] perform masked-language model (MLM) pre-training with multilingual text-only data for TTS. Considering the abundant resources of textual data in the real world, which are usually more readily accessible compared with audio, it is worthwhile to investigate the utilization of language models pre-trained on large text resources for phoneme representations in multilingual synthesis tasks.
@@
To synthesize speech with a target speaker’s timbre, end-to-end models usually learn a speaker embedding during training that enables speaker selection at inference time. Furthermore, several of these models enable zero-shot speaker synthesis using a speaker vector extracted from a short audio clip[13, 14]. Despite recent advances, the similarity gap between seen and unseen speakers remains an open research question, and training these models still requires a significant number of speakers, posing challenges in developing high-quality models for low-resource languages[8]. Furthermore, in real-world data, most of the attributes that represent speech are interrelated and difficult to disentangle. To control a multilingual and multispeaker system, it is crucial to disentangle speaker and language[6]. Efforts using adversarial learning or consistency loss [6, 8] have been made to mitigate the performance degradation resulting from this entanglement.
@@
However, these state-of-the-art multilingual and multispeaker systems mainly rely on a large amount of training data, which is not available for every language. The intermediate features used in these studies are often Mel spectrograms, which have a high correlation in time and frequency, making it difficult to disentangle speaker-dependent information. Fortunately, self-supervised learning (SSL) speech representations [15, 16, 17] have been shown to be useful for speech processing tasks, such as speech recognition [18, 19], speech reconstruction [20], and voice conversion [21]. The learned representations are used as input for a supervised model, which is often fine-tuned to improve task performance or reduce the labeled data required. Recently, several TTS models[22, 23] have started using discrete vector-quantized speech representations as intermediate features instead of traditional Mel spectrograms for prediction. As a result, the quantized output has less speaker-dependent information than the Mel spectrograms[23]. While a designed fine-tuning protocol for the learned representations indeed improves automatic speech recognition (ASR) performance [24], in low-resource settings, relatively little attention has been paid to the study of TTS models using learned SSL representations.
Recently, large-scale TTS systems [22, 25, 26, 27] that leverage data-driven representations, i.e., either discrete tokens or continuous vectors from an auto-encoder [28, 29] have been widely adopted for zero-shot speech synthesis. Several of these models [22, 25] employ an autoregressive architecture to generate discrete tokens one by one, like a language model. These autoregressive models suffer from a slow inference speed, unstable prosody, and word skipping/repeating issues. To overcome this limitation, non-autoregressive large-scale TTS systems, such as NaturalSpeech 2/3 [30, 31] and HierSpeech++ [32], have been investigated. However, all the aforementioned large-scale TTS systems heavily rely on abundant data and primarily focus on resource-rich languages rather than low-resource ones.
@@
In this paper, we propose a unified zero-shot multilingual multispeaker TTS framework called ZMM-TTS that leverages discrete speech representations from a multilingual speech-based SSL model. ZMM-TTS consists of a text-to-discrete speech representations module (txt2vec) and a representations-to-waveform module (vec2wav). Regarding the txt2vec module, we adopt a standard end-to-end TTS architecture comprising token embeddings, an encoder, and a decoder to predict discrete representations. To utilize the strong cross-lingual transferability of multilingual language models, we adopt a pre-trained large-scale multilingual language model for phoneme representations. For the vec2wav module, an additional multi-stage and multi-head vector quantization (VQ) model was adopted to preserve the discrete information at different time resolutions.
@@
  • We propose ZMM-TTS, a multilingual and multispeaker framework using discrete audio representations from a large-scale multilingual pre-trained self-supervised model as an intermediate representation to replace the widely-used Mel spectrograms.
  • We investigate the impact of various input representations on multilingual synthesis tasks that use SSL discrete representations. Our research combines phoneme representations from a pre-trained text-based multilingual language model and speech-based SSL representations in a speech synthesis task.
@@
  • We also verify the effectiveness of the ZMM-TTS in low-resource and zero-shot scenarios on two hypothetically low-resource languages. “Zero-shot” refers to the capability of synthesizing speech in a low-resource language without training or fine-tuning the model using the data from that language. This observation is quite promising, as it indicates that our proposed approach can synthesize intelligible audio with high speaker similarity, even when no specific training data are available for the previously unseen language. Note that “unseen” in this context refers to the absence of dedicated training data for ZMM-TTS, while the unseen language is still included in the training data of both the pre-trained language model and self-supervised model.
We encourage the reader to listen to our samples, which can be found at https://gongchenghhu.github.io/TASLP-demo/. The source code has been released on https://github.com/nii-yamagishilab/ZMM-TTS.
The remainder of this paper is organized as follows. Section II provides a comprehensive review of related work in the field. Section III presents a detailed description of the proposed approach. The experimental setup is described in Section IV. In Section V, we evaluate and analyze the proposed system using six languages, in particular, the quality of unseen speakers’ voices. Section VI evaluates and analyzes the proposed system, assuming two unseen low-resource languages. Finally, Section VII concludes the paper, summarizing the findings and highlighting future research directions. In Appendix A, we show a comparison with the latest large-scale models that utilize larger amounts of data.
@@ II RELATED WORK
This section reviews related work about zero-shot and multilingual speech synthesis. We also review related studies on speech synthesis with the aid of SSL representations, especially in the multilingual cases.
@@ II-A1 Zero-shot multispeaker TTS
The concept of zero-shot multispeaker TTS was initially introduced in [33] and the main idea is to employ embeddings from an external speaker encoder as a reference signal. Subsequent research by [14] aimed to reduce the quality gap between seen and unseen speakers by utilizing more informative embeddings. [34] further explored the utilization of attentive speaker embeddings for the purpose of encoding general speaking styles, while [35] proposed a speaker-conditional architecture and investigated a flow-based decoder capable of generating unseen speaker voices. Despite these efforts, the task of zero-shot multispeaker TTS has yet to be fully solved. Moreover, capturing a wide range of voice properties requires a substantial amount of high-quality data from numerous varied speakers.
Regarding multispeaker multilingual TTS, a primary idea is to train the decoder by incorporating learnable language and speaker embeddings[6]. However, the majority of speech attributes are interrelated and difficult to disentangle through embedding tables alone. Attempts have been made to mitigate the decline in multi-language multispeaker performance caused by speaker and language entanglement, especially in cross-lingual scenarios. [6] incorporated adversarial domain training, enabling them to transfer different voices between languages. [36] proposed the use of mutual information minimization to keep speaker consistency in cross-lingual synthesis. [37] utilized joint training incorporating a speaker classifier to enhance speaker similarity.
@@ II-A2 System structure
Previous multilingual TTS models [6, 7] are primarily built using Tacotron [1, 3]. When synthesizing speech, Tacotron-based models often repeat or skip words due to an autoregressive approach that uses attention to align input text and target speech. Instead of autoregressive iteration, several non-autoregressive [38, 39, 2, 40] models adopt Transformer encoder and decoder blocks to generate Mel spectrograms in parallel. For example, [41] create a multilingual speech synthesis system including a Transformer-based acoustic predictor and a WaveNet [42] neural vocoder. [43] presented a multilingual, multiaccented, multispeaker speech synthesis model based on RADTTS [44] which is a parallel flow-based generative model. Recently, several fully end-to-end multilingual speech synthesis systems [45, 8] were built on VITS[46], which could produce natural-sounding synthesized speech. Moreover, VITS requires training a single model, so there is no need for an independent vocoder. Among these fully end-to-end systems, YourTTS [8], which aims at zero-shot multispeaker TTS and zero-shot voice conversion, presented promising results for a few language combinations. However, it had limited success in transferring to languages with few speakers and used a curriculum learning approach, making training cumbersome.
@@ II-B1 Characters or Phonemes
Traditionally, end-to-end TTS models have utilized input representations in the form of characters or phonemes[1, 2]. Employing each character as the default input for end-to-end TTS models requires the model to implicitly learn how to convert graphemes to phonemes as part of the synthesis task. Incorporating phoneme input simplifies the TTS task, eliminating the need for the model to learn complicated grapheme-to-phoneme rules for languages such as English. Furthermore, language-independent pronunciation information can be obtained through language expertise, such as making a unified set of all the languages in a multilingual model, often derived from the International Phonetic Alphabet (IPA) [5, 47].
II-B2 Pre-trained representations
Over the past decade, self-supervised pre-training on extensive text corpora, employing language model (LM) or MLM objectives, has demonstrated remarkable success in the field of TTS[9, 10, 11]. Toward multilingual tasks, [12] proposed XPhoneBERT which is the first pre-trained multilingual model for phoneme representations for TTS. Based on the monolingual experiment results, using XPhoneBERT as a phoneme encoder for input has been shown to significantly enhance the performance of a powerful neural TTS model in terms of naturalness and prosody. It also helps to produce high-quality speech using low-resource training data. [5, 48] conducted pre-training on a multilingual text-only dataset using an MLM. They then trained this model in a supervised manner on paired text and audio data, while keeping the language-aware embedding layer frozen. This approach enables the model to synthesize languages that are not present in the paired data but exist in the text-only data.
Alternative approaches adopt UTF-8 bytes [49] or phonological features (PFs) [50, 51] as input representations. UTF-8 byte representations encode typographic rather than phonological information. The model does not learn unseen byte combinations, requiring retraining when enrolling in a new language. For PFs, IPA symbols can be converted to PFs in accordance with certain aspects of their articulation, as implemented in Epitran [52]. Although this conversion is based on an explicit database of IPA symbols and diacritics, it can be considered universally applicable to any language for which IPA transcription is available. However, there are no pre-trained phoneme-level encoders similar to XPhoneBERT currently available. The effect of different input representations is still not clear in multilingual speech synthesis with SSL representations. Taking into account the complexity of the model and the support for new languages and low-resource scenarios, we evaluated the effects of using characters, IPA, and XPhoneBERT as input for multilingual TTS.
@@ II-C SSL Representations for speech synthesis
Well-designed proxy tasks enable SSL to explore the characteristics of unlabeled data and enhance the model’s representation ability. The derived representations of the SSL model have useful information for many downstream tasks[53, 54, 55, 56, 57]. Wav2vec 2.0 [15] is the most popular pre-trained model that uses SSL on a large amount of unlabeled speech data, and it has demonstrated strong representation abilities in ASR tasks and achieved outstanding results. A multilingual wav2vec 2.0, denoted as XLSR-53, was also proposed in [58]. The results showed a significant improvement over monolingual wav2vec 2.0, particularly in languages with limited resources.
For SSL in monolingual TTS downstream tasks, [54] introduced WavThruVec, a two-stage architecture that employed high-dimensional wav2vec 2.0 embeddings as an intermediate speech representation. The paper [59] proposed VQTTS, which uses the vec2wav vocoder to train a non-autoregressive model that maps text to discrete tokens. MQ-TTS [60] proposed using multiple codebooks to quantize intermediate features on real-world spontaneous speech. Using wav2vec 2.0 acoustic representations of synthesized audio, [57] improved prosody modeling by incorporating prosodic characteristics from neighboring utterances.
@@
For multilingual tasks, [23] demonstrated that the VQ features in wav2vec 2.0 contain significantly less speaker information than the Mel spectrograms, and proposed DSE-TTS, a cross-lingual TTS that encodes timbre-related and linguistic-related speaker information separately. In the paper [61], the author proposed extracting unsupervised phonetic representations (UPR) from wav2vec 2.0 for multilingual speech synthesis when target language pronunciation dictionaries are unavailable. In the work of [62], the author built a TTS system for the Scottish Gaelic language based on self-supervised discrete acoustic unit sequences. Scaling TTS to many languages was proposed in [63] by leveraging massively multilingual joint speech and text representation learning.
@@
However, integrating SSL-based representations into TTS in low-resource settings is still a novel concept, and it is unclear which type of input representation is most suitable for TTS. Furthermore, these methods mostly use monolingual speech representations rather than multilingual speech representations, which may benefit low-resource scenarios.
II-D Large-scale TTS systems
Recently, considerable progress has been made in zero-shot TTS by scaling up corpus and model sizes. These large-scale TTS systems [22, 25, 26, 27] usually quantize the continuous speech waveform into discrete tokens through neural audio codec models and model these tokens with autoregressive LMs. For example, VALL-E [22] was the first neural codec language model for speech synthesis utilizing a discrete audio unit and language models. In addition to autoregressive architectures, several large-scale models [30, 31, 64, 32, 65] adopt non-autoregressive architectures to improve model generation speed and robustness. While several large-scale models [66, 27, 67], such as VALL-E-X [66], are multilingual and have been trained on extensive corpora in both Chinese and English, current large-scale models are more focused on generalization to unseen speakers rather than on adapting to unseen languages. These large-scale models perform well only in resource-rich languages while lacking application in low-resource languages.
The differences between our ZMM-TTS system and previous large-scale models are highlighted in Table LABEL:tab:tts_comparison. Compared with previous large-scale models, ZMM-TTS is designed to have the language adaptation ability for low-resource and unseen languages. Typically, previous models only validated one language or a few high-resource languages. In contrast, the ZMM-TTS model aims to support not only unseen speakers but also unseen languages.
TABLE I: Comparing ZMM-TTS with previous large-scale TTS models on task capabilities.
Systems Languages Few-shot Zero-shot
[25, 22, 31, 65, 64, 32, 26] Single Validated only on one language Unseen speakers
[66, 27, 67] Multiple Validated only on high-resource languages Unseen speakers
ZMM-TTS Multiple Adaptable to multiple low-resource languages Unseen speakers and languages